OpenStudy (anonymous):

Medal + Fan who ever gives me the correct answer first! What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2? y =___ x +___

OpenStudy (anonymous):

@Reaper534

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

ok thank you can i tag you again if I need help?

OpenStudy (anonymous):

yeah wait were not done

OpenStudy (anonymous):

oh ok :)

OpenStudy (anonymous):

(8,12) and slope -2 has to go in to Y=mx+b so its 12=-2(8) + b

OpenStudy (anonymous):

so 28?

OpenStudy (anonymous):

for B

OpenStudy (anonymous):

12=-2(8) + b no we slove for B 12=-16 +b +16 +16 28=b so the awnser is y=-2x+28

OpenStudy (anonymous):

ok Thank you!!!!!

OpenStudy (anonymous):

I was doing it wrong haha i Haven't done math in a while

OpenStudy (anonymous):

oh were you XD its okay lol

OpenStudy (anonymous):

haha yeah well link me if u need help

OpenStudy (anonymous):

I'll tag you in another q' if I need it :)

OpenStudy (anonymous):

if thats alright

OpenStudy (anonymous):

Thank you for the medal btw!